Solve the following system of equations by substitution and select the correct answer below:
4x-6y=28
8x-4y=24
a) x=1, y=4
b) x=-1, y=4
c) x=1, y=-4
d) x=-1, y=-4

Answer by Cromlix(4381)

Hi there,
4x - 6y = 28......(1)
8x - 4y = 24......(2)
........
Using Equation (2)
8x - 4y = 24
8x = 24 + 4y
x = 3 + 1/2y....(2)
.......
Substituting x = 3 + 1/2y
into Equation (1)
4(3 + 1/2y) - 6y = 28
12 + 2y - 6y = 28
Collect like terms:
2y - 6y = 28 - 12
-4y = 16
y = -4
Substitute y = -4 into
Equation (1)
4x - 6y = 28
4x - 6(-4) = 28
4x + 24 = 28
4x = 28 - 24
4x = 4
x = 1
x = 1; y = -4
Choice (c)
Hope this helps :-)